for observing newton's ring b/w with two different o media b/w glass surface .nth ring have daimeter 11:9 claculate ration of refractive index of two medium
1
Expert's answer
2018-09-25T10:19:08-0400
For the dark ring system:
(μd_n^2)/4R=nλ
μ_1 〖d_n^2〗_1=μ_2 〖d_n^2〗_2
The ratio of refractive index of two medium:
μ_1/μ_2 =(d_n2/d_n1 )^2=(9/11)^2=81/121.